OpenStudy (anonymous):

Can some please help me solve (reduce) this equation so that I can calculate C? |x|=e^[[ln(|V-F|) ]^((Q/J) )+C]

OpenStudy (across):

\[|x|=e^{\ln|V-F|^{\frac{Q}{J}}+C}?\]

OpenStudy (anonymous):

|x|=e^[[ln(|V-F|) ]^((Q/J) )+C] lxl=e^[ln lV-Fl ^(Q/J)] * e^C lxl / e^[ln lV-Fl ^(Q/J)] = e^C ln (lxl / e^[ln lV-Fl ^(Q/J)])= C

OpenStudy (jamesj):

In the across-jamesj school of thought, we try not to give the answer, but instead try to lead people to it. For example, here, once, we had confirmed the form of the equation, we might have asked: What is c in this equation y = e^(a+c) ==== But anyway ...